OpenStudy (anonymous):

Betty bought a package of meat that weighed 2 1/2 pounds and another package that weighed 1 3/4 pounds. How much meat did she buy all together?

OpenStudy (kropot72):

Put the mixed numbers into fractions as follows: \[2\frac{1}{2}=\frac{5}{2}\] \[1\frac{3}{4}=\frac{7}{4}\] Then add the fractions: \[\frac{5}{2}+\frac{7}{4}=\frac{10+7}{4}=?\] Can you solve it now?